OpenStudy (anonymous):

Can someone integrate this? (x^3)(e^x)

OpenStudy (anonymous):

do u know by-parts formula?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Do you need it? Because I've tried doing that and it just seemed to give me a giantic answer that didn't seem right.

OpenStudy (anonymous):

u have to select f and g out of the two if u did this correctly u wont stuck at all f=e^x g=x^3 now try it

OpenStudy (anonymous):

Okay, I'll try that. Thanks : )

OpenStudy (anonymous):

m i right , hmm m confused a bit , can u write the formula for sake of comfirmation

OpenStudy (anonymous):

The by parts formula? uv - v du

OpenStudy (anonymous):

oops - here I'll try again to type that

OpenStudy (anonymous):

nopes integration by parts formula

OpenStudy (anonymous):